MadisonMarvel:

The area of a square room can be expressed as 196x^12y^10z^6 square feet . Hajime is installing a wallpaper border along the perimeter of the room. What is the perimeter?

jhonyy9:

area = length time width area square = side time side = (side)^2 area = 196x^12y^10z^6 (side)^2 = 196x^12y^10z^6 side = sqrt(196x^12y^10z^6) side = (196x^12y^10z^6)^(1/2) side = 14x^(12/2)y^(10/2)z^(6/2) side = 14x^6y^5z^3 hope helped understandably easy @supie